1)Suppose that thepopulation variance of the amount of cooldrink supplied by avending machine is known to be 115𝑚𝑙2
i)The machine was activated 61times, and the mean amount of cooldrink supplied on each occasion was 𝑋̅. If the upper confidence limit for a 99% confidence level is 188.20. What is the value of the sample mean 𝑋̅?
ii)What sample size is requiredif the estimateis requiredto be within 1𝑚𝑙2of the true value with probability 0.95?

Answer by Boreal(15235)

sd=sqrt(115)=10.72 ml
the half-interval for 99%CI is z*sigma/sqrt(n)=2.576*10.72/sqrt(61)=3.54
therefore the mean is 188.20-3.54 or 184.66 ml
to be within 1 ml (not sure what the 2 means there); 115 ml^2 are appropriate units for variance, but ml is appropriate for the answer.
1.96*10.72/sqrt(n)=1; sqrt (n)=21.0112, n=441.47 or 442
sample size is 442.